How many electrons will be present in the sub-shells having ms value of +1/2 for n = 5 ?
 

Grade:11

1 Answers

Arun
6 years ago
 
  • There are 4 sub shells associated with shell n =4 ; because the Azimuthal quantum no., representing the sub-shells is from 0 to (n-1), i.e , 0,1,2...., (n-1).
  • As there are 4 sub shells, those will be :- s,p,d,f  giving a total of = 1+3+5+7= 16 orbitals and 1 electrons for each orbital with ms= -1/2 and +1/2 each.. Thus there will be 16 such electrons.